40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2030     Entry time: Thu Oct 1 03:12:56 2009
Author: rob 
Type: Update 
Category: Locking 
Subject: some progress 

Good progress in IFO locking tonight, with the arm powers reaching about half the full resonant maximum. 

Still to do is check out some weirdness with the OMC DAC, fix the wireless network, and look at c1susvme2 timing.

ELOG V3.1.3-